OpenStudy (anonymous):

Solve tan^(4)+2tan^(2)-3=0 for 0°≤θ<360°

OpenStudy (anonymous):

is it \[\tan 4 ??\]

OpenStudy (anonymous):

or \[\tan ^{4}x\]?

OpenStudy (anonymous):

latter one

OpenStudy (anonymous):

ok,take \[\tan^4x=y^2\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

y^(2)+2y-3=0

OpenStudy (anonymous):

wait a minute, i lost my calculator...

OpenStudy (anonymous):

then \[\tan^2x=y\] and the equation is \[y^2+2y-3=0\]

OpenStudy (anonymous):

y=1/-3

OpenStudy (anonymous):

tanx=1 x=45 tanx=-3 x=.....

OpenStudy (anonymous):

yep,so x=\[\tan^-1 (-3)\] use calculator.

OpenStudy (anonymous):

i lost my calculator..

OpenStudy (anonymous):

-71.565 Degrees -1.249 Radians

OpenStudy (anonymous):

:) it is -71.565 Degrees ,but \[\Theta>0\] so it will be 180 degrees-71.565 Degrees

OpenStudy (anonymous):

180-(-72.565)=252.656?

OpenStudy (anonymous):

also (360-71.565 )Degrees

OpenStudy (anonymous):

no,180+(-72.565) degree

OpenStudy (anonymous):

these are the choices:: A. 45, 135 only B. 45, 225 only C. 45, 60, 225, 240 D. 45,120,225,300 E. 45,135,225,315

OpenStudy (anonymous):

i think there is some problems with the answers.. cause , when \[\tan \alpha=\tan \beta \] \[\alpha=n \pi+\]

OpenStudy (anonymous):

beta

hartnn (hartnn):

hey, tan x is not -3 tan^2 x = -3

hartnn (hartnn):

so, discard that tan^2 x = 1 tan x = 1 , tan x = -1

OpenStudy (anonymous):

ohh,yeah yeah @hartnn is correct

OpenStudy (anonymous):

ohh, yes....

OpenStudy (anonymous):

45 and 180-(-45)=225

hartnn (hartnn):

you know why tan^2 x = -3 is discarded ??

OpenStudy (anonymous):

cause it will be imaginary

OpenStudy (anonymous):

ohh, imaginary number?

hartnn (hartnn):

yes, \(\tan x = \sqrt{-3}= i\sqrt{3}\) so no real solution

OpenStudy (anonymous):

okay...

hartnn (hartnn):

so what all angles you got finally ?

OpenStudy (anonymous):

45,225

hartnn (hartnn):

thats only for tan x = 1 what about tan x = -1 ?

hartnn (hartnn):

or, did you get why we need to also solve tan x = -1 ?

OpenStudy (anonymous):

135

OpenStudy (anonymous):

i know

OpenStudy (anonymous):

315

hartnn (hartnn):

135 and ?

OpenStudy (anonymous):

45,135,225,315 ?

hartnn (hartnn):

\(\huge \checkmark \)

OpenStudy (anonymous):

thanks

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

and i want to ask how to many how many roots are theere in a function

OpenStudy (anonymous):

i mean....like this: For 0°≤θ≤360°, how many roots does the equation sinx(cosx+2)=0 have?

hartnn (hartnn):

ok, sinx(cosx+2)=0 sin x = 0 (how many values of x in 0 to 360 ??) cos x +2 =0 cos x = -2 (how many values of x in 0 to 360 ??)

OpenStudy (anonymous):

3 for sinx=0

hartnn (hartnn):

are 0 and 360 considered different ?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

0, 180, 360

OpenStudy (anonymous):

i want to ask how can i press sin-1 in this website ? http://web2.0calc.com/

hartnn (hartnn):

i mean 0 and 360 are same angle,